### v1

# trans-subtitle [https://orcid.org/0000-0003-0875-8345] [https://orcid.org/0000-0002-6003-3168] [https://orcid.org/0009-0000-1789-4117] [https://orcid.org/0009-0000-8135-7260] [https://orcid.org/0000-0003-4632-5743] [https://orcid.org/0009-0000-5184-2695] [https://orcid.org/0000-0003-1324-1412]

Abstract
[ABSTRACT] Deploying nitrogen vacancy (NV) centers in diamond as nanoscale quantum sensors for condensed matter and materials physics requires placing the NV centers close to the sensing target. One solution is to fabricate diamond nanostructures and integrate them with materials and devices. However, diamond etching and ion milling can introduce subsurface damage and surface defects that degrade the charge stability and spin coherence of NV centers near the surface. Here we report a procedure for fabricating low-damage nanostructured diamond membranes, and we show that this fabrication scheme preserves the optical and spin properties of state-of-the-art shallow NV center quantum sensors, within nanometers of the diamond surface, while providing significant photonic enhancement. Furthermore, we demonstrate a pick-and-place transfer method, which enables integration with diverse sensing targets.
©This is an open access article under the terms of the Creative Commons Attribution-NonCommercial License, which permits use, distribution and reproduction in any medium, provided the original work is properly cited and is not used for commercial purposes.
© 2024 The Author(s) Advanced Functional Materials published by Wiley Periodicals LLC on behalf of Advanced Functional Materials.
keywords:
quantum sensing | diamond | nanofabrication | nitrogen vacancy | photonics | membrane | nanobeam
journal: Advanced Functional Materialsvolume: 0footnotetext: *These authors contributed equally to this work.
1 Introduction
[…]
There have been several advances in recent years in diamond nanofabrication, including reactive ion etching to define membranes, 19, pillars20, 21, 14 and nanobeams, 22, 23, 24 and smart-cut techniques to fabricate diamond-on-insulator devices. 25, 26, 27 An alternative method for realizing free-standing nanostructures is a modified version of the single crystal reactive etching and metallization (SCREAM) process, which was first developed for silicon 28 and recently adapted to single crystal, monolithic diamond. 29, 30 In this process, anisotropic, top-down etching is followed by the creation of a sidewall hardmask and subsequent quasi-isotropic etching to undercut the desired structures. In the quasi-isotropic etching step, setting the RF bias to zero induces a highly diffuse, isotropic, plasma that etches different crystallographic directions of diamond at different rates. 31 This strategy has been used to fabricate microdisk resonators 32 and diamond nanobeams. 30, 33, 34, 35, 36
Here we apply the quasi-isotropic etch as part of a process to fabricate a membrane consisting of a dense array of nanobeams, with sacrificial tethers to the host substrate. To achieve a highly uniform undercut across microns of dense nanostructures, we design the device to reduce variations in the quasi-isotropic etch rate. We perform the etch at lower temperature and inductively coupled plasma reactive ion etching (ICP-RIE) power than in reference 32, reducing the etch rate and enabling control over the undercut geometry as well as the beam and tether thicknesses. In particular, the sacrificial tether geometry can be tuned to achieve robust suspension while still allowing mechanical separation. This enables device integration through pick-and-place transfer via micro-manipulator probes.
By simultaneously ensuring that the surfaces have low contamination, low subsurface damage, and high quality oxygen termination, we fabricate nanobeams that host NV centers within nanometers of the surface that have state-of-the-art nanoscale sensing performance. We show that the spin coherence time, charge state stability, and charge dynamics of shallow NV centers in these nanobeams are in line with "typical best" bulk substrate processing. Furthermore, the nanobeam form factor enables collection efficiency enhancement by up to a factor of seven when integrated with a sapphire target, in agreement with photonic simulations.
Figure 1: (a) Fabrication scheme showing method for fabricating freestanding undercut structures. Steps 1-4 comprise the anisotropic etch of the SiNx hardmask, the anisotropic etch of diamond, followed by conformal PECVD with SiNx, then the final anisotropic etch of SiNx in the trench. Step 5 shows isotropic etching of three different width of nanobeams, resulting in different geometry of the underside. (b) Tilted SEM image (left) of nanobeam membrane, with a higher magnification image (right) of an individual nanobeam (720 nm wide and 300 nm thick). (c) SEM image shows the design of the trellis, tethers and gaps between the frame and host sample. Thinning of the diamond on the outer trellis corner section can also been seen. (d) FIB cross-sections of three different gold-coated nanobeam widths fabricated in one device run. The high contrast beads on the top of the beams result from the gold coating. (e) Optical image of the host diamond sample showing many nanobeam membranes across a large area of the 4.5mm x 4.5mm host sample. (f) SEM image overlaid with a confocal photoluminescence scan of a 720 nm wide nanobeam membrane acquired with air objective (NA = 0.9) under 520 nm (400 µW) illumination. Bright spots in the unfabricated host crystal and the nanobeams are NV centers.
2 Results
2.1 Diamond Nanobeam Membrane Fabrication
The fabrication process follows the approach of Mitchell et al. 32, with several adaptations to create high-density nanobeam membranes compatible with pick-and-place device integration for quantum sensing applications, as described below.
We prepare the surface of a (100) oriented electronic grade diamond (Element Six), according to the process outlined in 18 and implant it with 15N (3 keV, 1e9 cm-2) followed by an 800C activation anneal in vacuum to create shallow NV centers. We then deposit 150 nm of silicon nitride (SiNx) at 250C using plasma-enhanced chemical vapor deposition (PE-CVD, Oxford PlasmaPro 100), and pattern an etch mask (Zeon ZEP 520a) with electron beam lithography (Elionix ELS-F125), using a layer of spin-on conductive polymer (DisChem, DisCharge H2O X4) to reduce charging effects. The SiNx hard mask is patterned with a highly anisotropic etch (Figure 1a) using inductively-coupled plasma reactive ion etching (ICP RIE, PlasmaTherm Takachi SLR, 14 sccm CHF3 and 10 sccm SF6 at 5 mTorr, 200 W RF and 60 W ICP power). The high selectivity of the nitride in oxygen plasmas enables very smooth fabricated structures (Figure 1b).
We first fabricate the membranes with variable nanobeam widths (120 - 720 nm) to study how nanobeam width and thickness influence the fabrication procedure. Based on finite-difference time-domain modeling (FDTD) (Figure S.7 and Section 2.4), we target $\approx$700 nm width and $\approx$350 nm thickness, which sets the dimensions for all the features in a given fabrication run. Due to the upward etch rate of the quasi-isotropic etch, we etch to a deeper depth than the desired thickness during the anisotropic diamond etch, balancing these two etch rates to achieve the final desired thickness. Etching in pure oxygen plasma (10 mTorr, 110 W ICP-RIE, 1000 W RF Bias) for 165 s removes 820 nm of diamond with near vertical side walls, no edge collapse, and minimal micromasking (Figure S.1). Following the anisotropic etch, we deposit another 150 nm of SiNx conformally over the etched, SiNx-capped diamond mesa. The residual SiNx hard mask is thinned during the anisotropic etch step but is still continuous. The hard mask on the bottom diamond surface between the mesas is selectively removed using another directional SiNx etch process, leaving a SiNx hard mask layer on the sidewalls and top surface. Crucially, the resulting encapsulation of the corners of the device is sufficient to protect the top edge of the membrane during the quasi-isotropic etch.
The critical step is the long, quasi-isotropic etch performed in the same ICP-RIE with the platen heated to 140 °C. After an initial RF strike, the etch proceeds for 20 hours with 1000 W ICP and 0 W RF at 20 mTorr, resulting in quasi-isotropic etching along multiple directions. The lateral etch rate is strongly related to the distance between structures, as the quasi-isotropic etch rate is highly aspect-ratio dependent, necessitating long etches to achieve full undercut. For a 1 µm spacing, we find that the quasi-isotropic etch etches $\approx$600 nm laterally in 20 hours, guaranteeing full undercut for the targeted nanobeam width range. This aspect-ratio dependent rate is much smaller than in previous reports for structures such as microdisks, e.g. 2000 nm in 10 hrs ,29 and of sparse density rectangular photonic crystal cavities, 36 which do not suffer from an aspect ratio throttled etch rate.
We use focused ion beam (FIB) milling to examine the cross section of nanobeams of different widths that are etched at the same time (Figure 1d). The surface texture arises from gold and iridium coating to mitigate charging during milling. The thinner nanobeams effectively etch for longer, resulting in their bottom profile progressing from faceted to flattened over etching time. The asymmetry in the nanobeam profiles is due to wafer edge effects, where the lateral etch rate from the quasi-isotropic etch can vary across the sample due to proximity of the nearest wafer edge.
Another key aspect enabling the pick-and-place operation is careful design of the outer frame (“trellis”) and attachment tethers. The double-layer trellis remains rigid, enabling transferable nanobeam frames with individual nanobeams as narrow as 100 nm and tens of micrometers long. The standardized trellis results in a high yield process, enabling consistent undercutting of thousands of membranes across a 4.5 x 4.5 mm2 sample(Figure 1e). The 1 µm gap between the frame and host crystal ensures that the etch rate of the nanobeams is uniform even at the edges of the pattern, and the trellis buffers against aspect-ratio dependent edge effects, particularly at the corners (Figure 1c) where there is significantly faster upwards etching during the quasi-isotropic step. The tethers are very narrow (~100 nm) and thin to ensure mechanical flexibility37, and we pattern a large number to build a mechanically reliable anchor. These tethers can be broken individually by micromanipulators while allowing the membranes to withstand harsh processing. Without this design, frames would collapse or delaminate, precluding scalable, high-yield fabrication.
After isotropic etching, the sample is immersed in 49% HF for 24 hours to strip the SiNx, followed by cleaning in a refluxing mixture of nitric, perchloric, and sulfuric acids. Microscope images show that the resulting etched structures are smooth, and that the fabrication process has a high yield across a diamond wafer (Figure 1b,c,e).
We use X-ray photoelectron spectroscopy to verify surface cleanliness prior to performing a thermal oxygen anneal at 450C according to the procedure described in 18, which ensures high-quality oxygen termination. We image the resulting NV centers in a confocal microscope (NA = 0.9) under 520 nm illumination. The confocal microscope image overlaid with a scanning electron microscopy (SEM) image of the devices shows bright, single NV center emission in the nanobeams (Figure 1f). Single NV centers are brighter in the nanobeams compared to the host diamond, while their density is similar.
Figure 2: (a) Hahn echo coherence time T2 as a function of NV center depth in nanobeams and the host diamond, plotted together with data from Ref 18 for shallow NV centers after the oxygen surface termination procedure including a 1200C and only a 800C anneal (as in the current work). Yellow dashed line shows a $T_{2}=kd^{n}+T_{2,0}$ fit to the 800C annealed data from Ref. 18 ($k=0.047\pm 0.04,n=2.35\pm 0.25,T_{2,0}=3.35\pm 3.87$). Grey shaded area marks the error margin for $n$ with fixed $k=0.047$ and $T_{2,0}=3.35$ showing that most of the data that we measure in the nanobeams falls within this scaling range. (b) T2 measured using an XY-N pulse sequence as a function of the number of pulses, N, (N = 1 is a Hahn echo experiment) for two NV centers in the nanobeams. Dashed line shows $\propto{N^{2/3}}$ dependence. (c) Spectral noise density extracted from a XY-8 decoherence decay using Equation S2 for the same NV centers as in (b). The expected peak position associated with the 13C Larmor frequency is marked with a vertical dashed line. (d) $T_{1,SQ}$ and $T_{1,DQ}$ as a function of NV center depth in nanobeams (purple) and unfabricated diamond (blue).
2.2 NV Center Spin Properties
We examine the spin properties of NV centers implanted within nanometers of the diamond surface in nanobeams to evaluate their utility as nanoscale quantum sensors, and we compare them to those in the unfabricated regions of the diamond. We measure the Hahn echo coherence time ($T_{2}$) as a function of depth from the surface for several NV centers in the nanobeams and in the unfabricated diamond (Figure 2a). The depth is measured using proton nuclear magnetic resonance, as has been described previously. 38 The $T_{2}$ decreases as NV centers are brought closer to the diamond surface, as has been reported in the literature. 18, 39 The observed depth dependence of $T_{2}$ for NV centers in nanobeams is within the distribution of NV centers in the unfabricated regions, and is also consistent with state-of-the-art surface termination as described in the literature. 18 We note that the coherence times are slightly shorter than the best reported NV centers (Figure 2a, orange points), which resulted from a 1200C anneal that removes multivacancy clusters that form because of implantation damage, pointing to a potential direction for future improvement in the NV center performance in nanobeams.
To probe higher frequency noise, we measure the spin coherence with dynamical decoupling by XY-N pulse sequences. The coherence time $T_{2}$ increases with the number of pulses in the sequence, $N$, for two randomly selected NV centers (Figure 2b). We do not observe any saturation in $T_{2}$ up to 96 pulses, however the scaling deviates from the $N^{2/3}$} scaling expected for a slowly fluctuating noise bath,40 which can be attributed to the presence of broadband noise induced by defects at the diamond surface. Spectral noise decomposition using an XY-8 pulse sequence41 for the same two NV centers reveals a noise spectrum $S(\omega)$ consistent with a broadband noise bath and a pronounced peak corresponding to the 13C nuclear spin Larmor precession at 360 G (Figure 2c). These observations are consistent with previous characterization of the surface noise bath in state-of-the-art shallow NV centers. 18
Finally, we measure single-quantum spin lifetime ($T_{1,SQ}$) and double-quantum spin lifetime ($T_{1,DQ}$) to probe magnetic and electric noise near the NV center spin resonance frequency ($\approx$ 2.8 GHz)39, correspondingly, as a function of distance from the surface (Figure 2d). The spin lifetimes display a large distribution from 100 µs to several ms, with no clear depth dependence, and the distributions between NV centers in nanobeams and in the unfabricated region are similar. The $T_{1,DQ}$ values between 100 µs and 1 ms are similar to the results presented in 18, which confirms that the fabrication process does not contribute additional sources of electric noise from the diamond surface. At the same time, we observe several shallow NV centers with $T_{1,SQ}$ in the range of 200 µs to 1 ms compared to $T_{1,SQ}$ > 1 ms as reported in 18, which could point to slightly increased high-frequency magnetic noise at the surface.
[…]
2.3 NV Center Charge Properties
Surface defects, impurities, and subsurface traps induced by etching may also degrade the NV center charge state properties, 16, 42 resulting in lower charge initialization or poor spin contrast. We measure the steady state charge population of NV centers in the nanobeams and in the host diamond by recording the photon number histogram under weak ($\approx$3 µW) 594 nm orange illumination after conventional green initialization (Figure 3a). A double Poissonian distribution fit reveals the two charge states: NV- (higher photon counts) and NV0 (lower photon counts). We interrogate 116 NV centers in nanobeams of various widths and 148 NV centers in the host diamond and plot the resulting NV- population statistics in Figure 3b. The observed charge state distributions are very similar, with 50% and 45% of NV centers showing >60% NV- population in the nanobeams and the host diamond host, respectively.
It has been reported that NV spin contrast can be compromised due to increased photoionization rates in the presence of imperfect surface termination even when the steady state NV- population is normal. 16 Therefore, we measure the time-dependent photoluminescence under orange illumination of variable power and fit the individual photoluminescence decays to an exponential function to extract the corresponding ionization rates. The rates vary quadratically with power as $aP^{2}$, where $a$ is the ionization coefficient and $P$ is the power, as expected for two-photon photoionization and recombination (see Figure S.5). The statistics of the ionization coefficients $a$ for 61 NV centers in the nanobeams of various widths and for 85 NV centers in the host diamond are plotted in Figure 3c. We find the mean ionization rate coefficients to be 5.4 Hz/µW2 and 3.6 Hz/µW2 for nanobeams and for the host diamond, respectively, with $\approx$20% of the NV centers in nanobeams displaying $a$ > 10 Hz/µW2. This faster ionization can be attributed to the increased optical excitation efficiency in the nanobeams (discussed in more detail in Section 2.4).
To quantify the impact of this increased ionization rate on sensing performance, we measure the spin contrast by performing Rabi oscillations in a small magnetic field (25 G). The spin contrast is then calculated as $C=\frac{S_{0}-S_{1}}{S_{0}}$ where $S_{0}$ ($S_{1}$) is the signal measured after initializing the ground state spin of the NV center into a $m_{s}=0$ ($m_{s}=-1$) spin sublevel. The spin contrast distributions for 53 NV centers in the nanobeams of various widths and for 65 NV centers in the host diamond are shown in Figure 3d. 52% and 56% of NV centers in nanobeams and in the host diamond, respectively, are characterized by good spin contrast (>20%). We observe more NV centers with no spin contrast in the nanobeams (around 50% more), which we attribute to higher prevalence of NV centers with lower NV- population in this subset on the nanobeams (see Figure S.6). We conclude that the quality of the fabricated surfaces is compatible with adequate charge properties for nanoscale sensing.
2.4 Device Integration and Photonic Optimization
Figure 4: (a) Widefield image of a membrane mid-transfer, held electrostatically by two glass probes above the host diamond. (b) False-colored SEM of a 15um diamond membrane (red) fully transferred onto a sapphire substrate (grey). (c) Saturated photoluminescence intensity under 520 nm illumination for NV centers in the unfabricated host (left, blue) as compared to the nanobeams of various widths and thicknesses for suspended beams (left, purple) and beams transferred to a sapphire substrate (right, green). The solid red line represents the enhancement in collection efficiency averaged over different NV dipole orientations and lateral displacement. Dashed red lines represent the maximum and minimum collection efficiency enhancement from simulations for a given nanobeam geometry. (d) Saturation power for NV centers in the unfabricated host (left, blue), suspended beams (left, purple), and beams transferred to a sapphire substrate (right, green). Dashed black line in (c) and (d) marks the average values in the unfabricated host, while the center box plot line represents the median of the data.
It has been shown that pulled glass rods can be utilized for pick and place transfer of diamond cantilevers. 43 To transfer our membranes, we use two glass micromanipulator probes, fabricated via micropipette pulling and mounted on motorized stages in a customized 2D-material transfer microscope (HQ Graphene). Using these probes, we first break the membrane tethers, then pick up the membranes, which are electrostatically held to the pipette tip. (Figure 4a). Using both manipulators, membranes are moved, flipped, and placed onto the surface of a sapphire substrate (Figure 4b). By applying mechanical pressure with a micromanipulator we can ensure that membranes make good contact with the substrate. For a clean sapphire surface, transferred membranes have sufficient contact to remain in place after wet processing with refluxing nitric, perchloric and sulfuric acid, as well as with HF, piranha solution (2:1 sulfuric acid and hydrogen peroxide), solvents such as acetone, IPA, and water, as well as blow drying under nitrogen.
We measure the saturated NV center fluorescence rate and saturation power in nanobeams transferred onto sapphire and compare them to the NV centers in suspended nanobeams (Figure 4c). Both configurations experience an average collection efficiency enhancement over unfabricated host diamond, ~2x in air and ~3x on sapphire, with a large variation between the maximum and minimum possible values, up to a ~7x improvement. We also observe a commensurate reduction in the saturation power for NV centers within nanobeams, corresponding to an increased excitation efficiency (Figure 4d), corroborating measurements of increased ionization rates in the nanobeams (Figure 3c).
FDTD simulations (Lumerical) are used to model the collection efficiency as a function of the nanobeam geometry, NV position, dipole orientation, and index of refraction of the environment. We approximate the nanobeam cross sections as perfect rectangles. Nitrogen implantation introduces stochastic variation in depth and lateral position of NV centers with respect to the nanobeam. We assume a 6 nm implantation depth without straggle, given a minor dependence of the collection efficiency on NV depth (Figure S.8). There is a strong dependence of the collection efficiency on the lateral position of the NV center due to the proximity of the diamond sidewall-air interface. Furthermore, different dipole orientations, namely, the pair of dipoles oriented toward the objective, and the pair oriented away from the objective in (100) diamond, are brighter and darker, respectively. Consequently, the possible range of enhancement is large, with some points in this space experiencing no enhancement and some reaching up to a 7x improvement in collection efficiency. The maximum and minimum collection efficiency for a given nanobeam geometry is shown as a red-dashed line (Figure 4c), highlighting this spread. A detailed analysis of this variation can be found in the Supporting Information (Section S.4). We perform this parameter sweep for two different sensing configurations: suspended nanobeams and nanobeams heterogeneously integrated (implanted-side down) onto a sapphire substrate. When flipped, the NV center position is close to the substrate, changing both the effective refractive index of the environment, as well as its relative position with respect to the objective and the nanobeam. This position benefits from total internal reflection at the nanobeam sidewalls, waveguiding the emission and enhancing the overall collection efficiency compared to the suspended case. We observe good agreement between the simulated and the observed enhancement in saturated fluorescence (Figure 4c). While we do not measure the precise lateral displacement beyond the optical diffraction limit, we observe variation in the collection efficiency that is consistent with the simulated distribution.
3 Conclusions
We establish nanostructured diamond membranes that host state-of-the-art shallow NV centers as a practical platform for quantum sensing experiments. The membrane fabrication allows for long NV center spin coherence and stable charge dynamics, while structuring the membrane into nanobeams enables enhanced photon collection. We have demonstrated a device geometry that is compatible with polymer-free, pick-and-place mechanical transfer, which can be readily applied to sensing in a variety of superconducting, magnetic, and other condensed matter systems. In particular, this sensing geometry aids in the interrogation of optically opaque materials, which is a particularly challenging task for single-crystal diamond sensors. Additionally, enhanced collection efficiency and reduction in required excitation power improves measurement throughput and reduces optical illumination of the target material. Future potential extensions of these capabilities include patterned implantation of NV centers using a nanopatterned mask 44 or nanoimplantation techniques, 45 which would provide more reliable collection efficiency enhancement through control of the NV center lateral position, particularly with preferential dipole orientation alignment in (111)-oriented substrates. 46 This procedure could be extended to the fabrication of low-dimensional structures such as diamond nanosheets and nanowires, opening up a novel avenue for studying carrier47 and heat48 transport in diamond.
[…]
*Supporting Information Supporting Information is available from the Wiley Online Library or from the author
[…]
*Conflict of Interest
The authors declare no conflict of interest.
[…]
*Acknowledgments This work was primarily supported by the Gordon and Betty Moore Foundation (grant GBMF12237, DOI 10.37807) and by the National Science Foundation under Grant No. OMA-2326767. Low-damage diamond etching was primarily supported by the U.S. Department of Energy, Office of Science, Office of Fusion Energy Sciences, under Award Number DE-SC0024537.
We thank Alex Abulnaga for help with the initial FDTD code we used in this work. The authors acknowledge the use of Princeton’s Imaging and Analysis Center (IAC), which is partially supported by the Princeton Center for Complex Materials (PCCM), a National Science Foundation (NSF) Materials Research Science and Engineering Center (MRSEC; DMR-2011750) and John J. Schreiber for assistance in FIB and imaging. The devices in this work were fabricated at the Princeton Materials Institute Micro/Nanofabrication Center and Quantum Device Nanofabrication Laboratory - Princeton University. We also thank HQ graphene for their help and support building the micromanipulator transfer microscope.
[…]
*Author contributions A.C.P-T. developed the design and fabrication process, and performed measurements. A.L. carried out the spin and charge measurements. S.K. assisted with fabrication development and with T.B.M. conducted the modeling, photonic simulations, and PL measurements. M.L. performed the FIB cross-sectional analysis and assisted with fabrication. K.-H.C. and M.L. built the confocal microscope used for PL measurements. N.P.dL. supervised the project. A.C.P-T, A.L, S.K and N.P.dL wrote the manuscript.
[…]
*Financial Disclosure
None reported.
[…]
*Conflicts of Interest
The authors declare no conflicts of interest.
[…]
*Data Availability
The data that support the findings of this study are available from the corresponding author upon reasonable request
[…]
*Keywords quantum sensing | diamond | nanofabrication | nitrogen vacancy | photonics | quantum sensor integration| membrane | nanobeam

### v2

@@
-trans-subtitle [https://orcid.org/0000-0003-0875-8345] [https://orcid.org/0000-0002-6003-3168] [https://orcid.org/0009-0000-1789-4117] [https://orcid.org/0009-0000-8135-7260] [https://orcid.org/0000-0003-4632-5743] [https://orcid.org/0009-0000-5184-2695] [https://orcid.org/0000-0003-1324-1412]
-Abstract
-[ABSTRACT] Deploying nitrogen vacancy (NV) centers in diamond as nanoscale quantum sensors for condensed matter and materials physics requires placing the NV centers close to the sensing target.
+High-quality nanostructured diamond membranes for nanoscale quantum sensing
+High-quality nanostructured diamond membranes for nanoscale quantum sensing
+Alexander C Pakpour-Tabrizi These authors contributed equally to this work.
+Princeton University, Department of Electrical and Computer Engineering, Princeton, NJ, USA    Artur Lozovoi These authors contributed equally to this work.
+Princeton University, Department of Electrical and Computer Engineering, Princeton, NJ, USA    Sean Karg These authors contributed equally to this work.
+Princeton University, Department of Electrical and Computer Engineering, Princeton, NJ, USA    Tecla Bottinelli Montandon University of Basel, Department of Physics, Basel, Switzerland    Melody Leung Princeton University, Department of Electrical and Computer Engineering, Princeton, NJ, USA    Kai-Hung Cheng Princeton University, Department of Electrical and Computer Engineering, Princeton, NJ, USA    Nathalie P.
+de Leon npdeleon@princeton.edu Princeton University, Department of Electrical and Computer Engineering, Princeton, NJ, USA
+(November 16, 2025)
+Deploying nitrogen vacancy (NV) centers in diamond as nanoscale quantum sensors for condensed matter and materials physics requires placing the NV centers close to the sensing target.
 One solution is to fabricate diamond nanostructures and integrate them with materials and devices.
@@
 Furthermore, we demonstrate a pick-and-place transfer method, which enables integration with diverse sensing targets.
-©This is an open access article under the terms of the Creative Commons Attribution-NonCommercial License, which permits use, distribution and reproduction in any medium, provided the original work is properly cited and is not used for commercial purposes.
-© 2024 The Author(s) Advanced Functional Materials published by Wiley Periodicals LLC on behalf of Advanced Functional Materials.
-keywords:
-quantum sensing | diamond | nanofabrication | nitrogen vacancy | photonics | membrane | nanobeam
-journal: Advanced Functional Materialsvolume: 0footnotetext: *These authors contributed equally to this work.
 1 Introduction
 […]
-There have been several advances in recent years in diamond nanofabrication, including reactive ion etching to define membranes, 19, pillars20, 21, 14 and nanobeams, 22, 23, 24 and smart-cut techniques to fabricate diamond-on-insulator devices.
-25, 26, 27 An alternative method for realizing free-standing nanostructures is a modified version of the single crystal reactive etching and metallization (SCREAM) process, which was first developed for silicon 28 and recently adapted to single crystal, monolithic diamond.
-29, 30 In this process, anisotropic, top-down etching is followed by the creation of a sidewall hardmask and subsequent quasi-isotropic etching to undercut the desired structures.
+There have been several advances in recent years in diamond nanofabrication, including reactive ion etching to define membranes, [corazza2025homogeneous], pillars[Wrachtrup_pillar, Maletinsky_AFMCryo, scheidegger2022scanning] and nanobeams, [Loncar_beams_fab, Yacoby_tribeams_chempot, Du_tribeams_Fe] and smart-cut techniques to fabricate diamond-on-insulator devices.
+[Membrane_2_High, guo_direct-bonded_2024, ding_high-q_2024] An alternative method for realizing free-standing nanostructures is a modified version of the single crystal reactive etching and metallization (SCREAM) process, which was first developed for silicon [SCREAM] and recently adapted to single crystal, monolithic diamond.
+[Khanaliloo_highQ, Barclay_waveguide] In this process, anisotropic, top-down etching is followed by the creation of a sidewall hardmask and subsequent quasi-isotropic etching to undercut the desired structures.
 In the quasi-isotropic etching step, setting the RF bias to zero induces a highly diffuse, isotropic, plasma that etches different crystallographic directions of diamond at different rates.
-31 This strategy has been used to fabricate microdisk resonators 32 and diamond nanobeams.
-30, 33, 34, 35, 36
+[Yacoby_isodeepdive] This strategy has been used to fabricate microdisk resonators [Barclay_300000] and diamond nanobeams.
+[Barclay_waveguide, TeonoFibreCoupled, hanson_waveguide, brevoord_large-range_2025, Englund_phc]
 Here we apply the quasi-isotropic etch as part of a process to fabricate a membrane consisting of a dense array of nanobeams, with sacrificial tethers to the host substrate.
 To achieve a highly uniform undercut across microns of dense nanostructures, we design the device to reduce variations in the quasi-isotropic etch rate.
-We perform the etch at lower temperature and inductively coupled plasma reactive ion etching (ICP-RIE) power than in reference 32, reducing the etch rate and enabling control over the undercut geometry as well as the beam and tether thicknesses.
+We perform the etch at lower temperature and inductively coupled plasma reactive ion etching (ICP-RIE) power than in reference [Barclay_300000], reducing the etch rate and enabling control over the undercut geometry as well as the beam and tether thicknesses.
 In particular, the sacrificial tether geometry can be tuned to achieve robust suspension while still allowing mechanical separation.
@@
 By simultaneously ensuring that the surfaces have low contamination, low subsurface damage, and high quality oxygen termination, we fabricate nanobeams that host NV centers within nanometers of the surface that have state-of-the-art nanoscale sensing performance.
-We show that the spin coherence time, charge state stability, and charge dynamics of shallow NV centers in these nanobeams are in line with "typical best" bulk substrate processing.
+We show that the spin coherence time, charge state stability, and charge dynamics of shallow NV centers in these nanobeams are in line with ”typical best” bulk substrate processing.
 Furthermore, the nanobeam form factor enables collection efficiency enhancement by up to a factor of seven when integrated with a sapphire target, in agreement with photonic simulations.
@@
 The fabrication process follows the approach of Mitchell et al.
-32, with several adaptations to create high-density nanobeam membranes compatible with pick-and-place device integration for quantum sensing applications, as described below.
-We prepare the surface of a (100) oriented electronic grade diamond (Element Six), according to the process outlined in 18 and implant it with 15N (3 keV, 1e9 cm-2) followed by an 800C activation anneal in vacuum to create shallow NV centers.
+[Barclay_300000], with several adaptations to create high-density nanobeam membranes compatible with pick-and-place device integration for quantum sensing applications, as described below.
+We prepare the surface of a (100) oriented electronic grade diamond (Element Six), according to the process outlined in [Sangtawesin_SurfNoise] and implant it with 15N (3 keV, 1e9 cm-2) followed by an 800C activation anneal in vacuum to create shallow NV centers.
 We then deposit 150 nm of silicon nitride (SiNx) at 250C using plasma-enhanced chemical vapor deposition (PE-CVD, Oxford PlasmaPro 100), and pattern an etch mask (Zeon ZEP 520a) with electron beam lithography (Elionix ELS-F125), using a layer of spin-on conductive polymer (DisChem, DisCharge H2O X4) to reduce charging effects.
@@
 We first fabricate the membranes with variable nanobeam widths (120 - 720 nm) to study how nanobeam width and thickness influence the fabrication procedure.
-Based on finite-difference time-domain modeling (FDTD) (Figure S.7 and Section 2.4), we target $\approx$700 nm width and $\approx$350 nm thickness, which sets the dimensions for all the features in a given fabrication run.
+Based on finite-difference time-domain modeling (FDTD) (Figure S.7 and Section S.4), we target $\approx$700 nm width and $\approx$350 nm thickness, which sets the dimensions for all the features in a given fabrication run.
 Due to the upward etch rate of the quasi-isotropic etch, we etch to a deeper depth than the desired thickness during the anisotropic diamond etch, balancing these two etch rates to achieve the final desired thickness.
@@
 This aspect-ratio dependent rate is much smaller than in previous reports for structures such as microdisks, e.g.
-2000 nm in 10 hrs ,29 and of sparse density rectangular photonic crystal cavities, 36 which do not suffer from an aspect ratio throttled etch rate.
+2000 nm in 10 hrs ,[Khanaliloo_highQ] and of sparse density rectangular photonic crystal cavities, [Englund_phc] which do not suffer from an aspect ratio throttled etch rate.
 We use focused ion beam (FIB) milling to examine the cross section of nanobeams of different widths that are etched at the same time (Figure 1d).
@@
 The 1 µm gap between the frame and host crystal ensures that the etch rate of the nanobeams is uniform even at the edges of the pattern, and the trellis buffers against aspect-ratio dependent edge effects, particularly at the corners (Figure 1c) where there is significantly faster upwards etching during the quasi-isotropic step.
-The tethers are very narrow (~100 nm) and thin to ensure mechanical flexibility37, and we pattern a large number to build a mechanically reliable anchor.
+The tethers are very narrow (~100 nm) and thin to ensure mechanical flexibility[Suresh_mechFlex], and we pattern a large number to build a mechanically reliable anchor.
 These tethers can be broken individually by micromanipulators while allowing the membranes to withstand harsh processing.
@@
 Microscope images show that the resulting etched structures are smooth, and that the fabrication process has a high yield across a diamond wafer (Figure 1b,c,e).
-We use X-ray photoelectron spectroscopy to verify surface cleanliness prior to performing a thermal oxygen anneal at 450C according to the procedure described in 18, which ensures high-quality oxygen termination.
+We use X-ray photoelectron spectroscopy to verify surface cleanliness prior to performing a thermal oxygen anneal at 450C according to the procedure described in [Sangtawesin_SurfNoise], which ensures high-quality oxygen termination.
 We image the resulting NV centers in a confocal microscope (NA = 0.9) under 520 nm illumination.
@@
 Single NV centers are brighter in the nanobeams compared to the host diamond, while their density is similar.
-Figure 2: (a) Hahn echo coherence time T2 as a function of NV center depth in nanobeams and the host diamond, plotted together with data from Ref 18 for shallow NV centers after the oxygen surface termination procedure including a 1200C and only a 800C anneal (as in the current work).
+Figure 2: (a) Hahn echo coherence time T2 as a function of NV center depth in nanobeams and the host diamond, plotted together with data from Ref [Sangtawesin_SurfNoise] for shallow NV centers after the oxygen surface termination procedure including a 1200C and only a 800C anneal (as in the current work).
 Yellow dashed line shows a $T_{2}=kd^{n}+T_{2,0}$ fit to the 800C annealed data from Ref.
-18 ($k=0.047\pm 0.04,n=2.35\pm 0.25,T_{2,0}=3.35\pm 3.87$).
+[Sangtawesin_SurfNoise] ($k=0.047\pm 0.04,n=2.35\pm 0.25,T_{2,0}=3.35\pm 3.87$).
 Grey shaded area marks the error margin for $n$ with fixed $k=0.047$ and $T_{2,0}=3.35$ showing that most of the data that we measure in the nanobeams falls within this scaling range.
@@
 The depth is measured using proton nuclear magnetic resonance, as has been described previously.
-38 The $T_{2}$ decreases as NV centers are brought closer to the diamond surface, as has been reported in the literature.
-18, 39 The observed depth dependence of $T_{2}$ for NV centers in nanobeams is within the distribution of NV centers in the unfabricated regions, and is also consistent with state-of-the-art surface termination as described in the literature.
-18 We note that the coherence times are slightly shorter than the best reported NV centers (Figure 2a, orange points), which resulted from a 1200C anneal that removes multivacancy clusters that form because of implantation damage, pointing to a potential direction for future improvement in the NV center performance in nanobeams.
+[pham2016nmr] The $T_{2}$ decreases as NV centers are brought closer to the diamond surface, as has been reported in the literature.
+[Sangtawesin_SurfNoise, myers2014probing] The observed depth dependence of $T_{2}$ for NV centers in nanobeams is within the distribution of NV centers in the unfabricated regions, and is also consistent with state-of-the-art surface termination as described in the literature.
+[Sangtawesin_SurfNoise] We note that the coherence times are slightly shorter than the best reported NV centers (Figure 2a, orange points), which resulted from a 1200C anneal that removes multivacancy clusters that form because of implantation damage, pointing to a potential direction for future improvement in the NV center performance in nanobeams.
 To probe higher frequency noise, we measure the spin coherence with dynamical decoupling by XY-N pulse sequences.
 The coherence time $T_{2}$ increases with the number of pulses in the sequence, $N$, for two randomly selected NV centers (Figure 2b).
-We do not observe any saturation in $T_{2}$ up to 96 pulses, however the scaling deviates from the $N^{2/3}$} scaling expected for a slowly fluctuating noise bath,40 which can be attributed to the presence of broadband noise induced by defects at the diamond surface.
-Spectral noise decomposition using an XY-8 pulse sequence41 for the same two NV centers reveals a noise spectrum $S(\omega)$ consistent with a broadband noise bath and a pronounced peak corresponding to the 13C nuclear spin Larmor precession at 360 G (Figure 2c).
+We do not observe any saturation in $T_{2}$ up to 96 pulses, however the scaling deviates from the $N^{2/3}$} scaling expected for a slowly fluctuating noise bath,[bar2012suppression] which can be attributed to the presence of broadband noise induced by defects at the diamond surface.
+Spectral noise decomposition using an XY-8 pulse sequence[romach2015spectroscopy] for the same two NV centers reveals a noise spectrum $S(\omega)$ consistent with a broadband noise bath and a pronounced peak corresponding to the 13C nuclear spin Larmor precession at 360 G (Figure 2c).
 These observations are consistent with previous characterization of the surface noise bath in state-of-the-art shallow NV centers.
-18
-Finally, we measure single-quantum spin lifetime ($T_{1,SQ}$) and double-quantum spin lifetime ($T_{1,DQ}$) to probe magnetic and electric noise near the NV center spin resonance frequency ($\approx$ 2.8 GHz)39, correspondingly, as a function of distance from the surface (Figure 2d).
+[Sangtawesin_SurfNoise]
+Finally, we measure single-quantum spin lifetime ($T_{1,SQ}$) and double-quantum spin lifetime ($T_{1,DQ}$) to probe magnetic and electric noise near the NV center spin resonance frequency ($\approx$ 2.8 GHz)[myers2014probing], correspondingly, as a function of distance from the surface (Figure 2d).
 The spin lifetimes display a large distribution from 100 µs to several ms, with no clear depth dependence, and the distributions between NV centers in nanobeams and in the unfabricated region are similar.
-The $T_{1,DQ}$ values between 100 µs and 1 ms are similar to the results presented in 18, which confirms that the fabrication process does not contribute additional sources of electric noise from the diamond surface.
-At the same time, we observe several shallow NV centers with $T_{1,SQ}$ in the range of 200 µs to 1 ms compared to $T_{1,SQ}$ > 1 ms as reported in 18, which could point to slightly increased high-frequency magnetic noise at the surface.
+The $T_{1,DQ}$ values between 100 µs and 1 ms are similar to the results presented in [Sangtawesin_SurfNoise], which confirms that the fabrication process does not contribute additional sources of electric noise from the diamond surface.
+At the same time, we observe several shallow NV centers with $T_{1,SQ}$ in the range of 200 µs to 1 ms compared to $T_{1,SQ}$ ¿ 1 ms as reported in [Sangtawesin_SurfNoise], which could point to slightly increased high-frequency magnetic noise at the surface.
 […]
 2.3 NV Center Charge Properties
-Surface defects, impurities, and subsurface traps induced by etching may also degrade the NV center charge state properties, 16, 42 resulting in lower charge initialization or poor spin contrast.
+Surface defects, impurities, and subsurface traps induced by etching may also degrade the NV center charge state properties, [yuan2020charge, dhomkar2018charge] resulting in lower charge initialization or poor spin contrast.
 We measure the steady state charge population of NV centers in the nanobeams and in the host diamond by recording the photon number histogram under weak ($\approx$3 µW) 594 nm orange illumination after conventional green initialization (Figure 3a).
@@
 We interrogate 116 NV centers in nanobeams of various widths and 148 NV centers in the host diamond and plot the resulting NV- population statistics in Figure 3b.
-The observed charge state distributions are very similar, with 50% and 45% of NV centers showing >60% NV- population in the nanobeams and the host diamond host, respectively.
+The observed charge state distributions are very similar, with 50% and 45% of NV centers showing ¿60% NV- population in the nanobeams and the host diamond host, respectively.
 It has been reported that NV spin contrast can be compromised due to increased photoionization rates in the presence of imperfect surface termination even when the steady state NV- population is normal.
-16 Therefore, we measure the time-dependent photoluminescence under orange illumination of variable power and fit the individual photoluminescence decays to an exponential function to extract the corresponding ionization rates.
+[yuan2020charge] Therefore, we measure the time-dependent photoluminescence under orange illumination of variable power and fit the individual photoluminescence decays to an exponential function to extract the corresponding ionization rates.
 The rates vary quadratically with power as $aP^{2}$, where $a$ is the ionization coefficient and $P$ is the power, as expected for two-photon photoionization and recombination (see Figure S.5).
 The statistics of the ionization coefficients $a$ for 61 NV centers in the nanobeams of various widths and for 85 NV centers in the host diamond are plotted in Figure 3c.
-We find the mean ionization rate coefficients to be 5.4 Hz/µW2 and 3.6 Hz/µW2 for nanobeams and for the host diamond, respectively, with $\approx$20% of the NV centers in nanobeams displaying $a$ > 10 Hz/µW2.
+We find the mean ionization rate coefficients to be 5.4 Hz/µW2 and 3.6 Hz/µW2 for nanobeams and for the host diamond, respectively, with $\approx$20% of the NV centers in nanobeams displaying $a$ ¿ 10 Hz/µW2.
 This faster ionization can be attributed to the increased optical excitation efficiency in the nanobeams (discussed in more detail in Section 2.4).
@@
 The spin contrast distributions for 53 NV centers in the nanobeams of various widths and for 65 NV centers in the host diamond are shown in Figure 3d.
-52% and 56% of NV centers in nanobeams and in the host diamond, respectively, are characterized by good spin contrast (>20%).
+52% and 56% of NV centers in nanobeams and in the host diamond, respectively, are characterized by good spin contrast (>>20%).
 We observe more NV centers with no spin contrast in the nanobeams (around 50% more), which we attribute to higher prevalence of NV centers with lower NV- population in this subset on the nanobeams (see Figure S.6).
@@
 It has been shown that pulled glass rods can be utilized for pick and place transfer of diamond cantilevers.
-43 To transfer our membranes, we use two glass micromanipulator probes, fabricated via micropipette pulling and mounted on motorized stages in a customized 2D-material transfer microscope (HQ Graphene).
+[Maletinsky_tran] To transfer our membranes, we use two glass micromanipulator probes, fabricated via micropipette pulling and mounted on motorized stages in a customized 2D-material transfer microscope (HQ Graphene).
 Using these probes, we first break the membrane tethers, then pick up the membranes, which are electrostatically held to the pipette tip.
@@
 Additionally, enhanced collection efficiency and reduction in required excitation power improves measurement throughput and reduces optical illumination of the target material.
-Future potential extensions of these capabilities include patterned implantation of NV centers using a nanopatterned mask 44 or nanoimplantation techniques, 45 which would provide more reliable collection efficiency enhancement through control of the NV center lateral position, particularly with preferential dipole orientation alignment in (111)-oriented substrates.
-46 This procedure could be extended to the fabrication of low-dimensional structures such as diamond nanosheets and nanowires, opening up a novel avenue for studying carrier47 and heat48 transport in diamond.
-[…]
-*Supporting Information Supporting Information is available from the Wiley Online Library or from the author
-[…]
-*Conflict of Interest
-The authors declare no conflict of interest.
-[…]
-*Acknowledgments This work was primarily supported by the Gordon and Betty Moore Foundation (grant GBMF12237, DOI 10.37807) and by the National Science Foundation under Grant No.
+Future potential extensions of these capabilities include patterned implantation of NV centers using a nanopatterned mask [implant_aligned] or nanoimplantation techniques, [FIB_implant] which would provide more reliable collection efficiency enhancement through control of the NV center lateral position, particularly with preferential dipole orientation alignment in (111)-oriented substrates.
+[osterkamp_engineering_2019] This procedure could be extended to the fabrication of low-dimensional structures such as diamond nanosheets and nanowires, opening up a novel avenue for studying carrier[lozovoi2021optical] and heat[li2012thermal] transport in diamond.
+Acknowledgements.
+This work was primarily supported by the Gordon and Betty Moore Foundation (grant GBMF12237, DOI 10.37807) and by the National Science Foundation under Grant No.
 OMA-2326767.
@@
 We also thank HQ graphene for their help and support building the micromanipulator transfer microscope.
-[…]
-*Author contributions A.C.P-T.
+A.C.P-T.
 developed the design and fabrication process, and performed measurements.
@@
 A.C.P-T, A.L, S.K and N.P.dL wrote the manuscript.
-[…]
-*Financial Disclosure
-None reported.
-[…]
-*Conflicts of Interest
-The authors declare no conflicts of interest.
-[…]
-*Data Availability
-The data that support the findings of this study are available from the corresponding author upon reasonable request
-[…]
-*Keywords quantum sensing | diamond | nanofabrication | nitrogen vacancy | photonics | quantum sensor integration| membrane | nanobeam
+Data availability: The data that support the findings of this study are available from the corresponding author upon reasonable request.